know that in 2006, there is 73% of adults used in internet. a random sample of 120 adults show that there are 95 of them used the internet. use 1% significance level and p-value method to test the claim that the current percentage of canadian adults who now use the internet has increased.
Answer by Boreal(15235)

will use 1-way
Ho: use of internet <=0.75
Ha: use of internet >0.75
alpha=0.05, p{reject Ho|Ho true)
test stat 1-sample proportion
z=(phat-po)/sqrt(p*(1-p)/n)
reject Ho for z>=2.32
z=(0.792-0.75)/sqrt (0.75*.25/120)
z=0.042/0.0395=1.06
fail to reject Ho at the 1% level
p-value is z>1.06 or 0.145